# On the arithmetic of polynomial ideals

### v1

@@ 3.1. Atoms constructed from sum-free sets of positive integers
We are now in the position to extend the results on the atomicity of $\mathfrak{b}_{i}(X,Y)$ and $\mathfrak{c}_{2i+1}(X,Y)$ to a larger family of ideals. Before proceeding, let us recall a classical definition from additive combinatorics: a nonempty subset $A\subseteq\mathbb{N}$ is sum-free if $A\cap(A+A)=\emptyset$. Easy examples of sum-free sets include any set of odd positive integers and any subset of $\{\lceil{n/2}\rceil,\dots,n\}$, where $n\in\mathbb{N}$ and $\lceil\ \rceil$ denotes the usual ceiling function. In fact, Green proved in [28] that these subsets account for “most” of the sum-free sets; more precisely, he proved that the number of sum-free subsets of $\{1,\dots,n\}$ is $O(2^{\nicefrac{{n}}{{2}}})$ for $n\in\mathbb{N}$, as predicted by the Cameron-Erdös Conjecture.
@@
For every $i\in\mathbb{N}^{+}$, the sets $\beta_{i}=\{i\}$ and $\delta_{2i+1}=\{2\llbracket 0,i\rrbracket+1\}$ are clearly sum-free. We then get from Lemma 3.2 an alternative proof of the fact (mentioned in the introduction and proved in [16]) that $\{0\}\cup\beta_{i}$ and $\{0\}\cup\delta_{2i+1}$ are atoms of $\mathcal{P}_{\rm fin}(\mathbb{N})$. The next theorem (the first of the main results of this section) provides a general method for constructing atoms in ${\mathcal{I}}(R)$ from sum-free subsets of $\mathbb{N}$ and yields, as a special case, the atomicity of $\mathfrak{b}_{i}(X,Y)=I_{\{0\}\cup\beta_{i}}$ and $\mathfrak{c}_{2i+1}(X,Y)=I_{\{0\}\cup\delta_{2i+1}}$.
@@ 3.2. A further family of atoms
For every $i\in\mathbb{N}_{\geq 2}$, the set $\delta_{2i}=\{1,2\llbracket 1,i\rrbracket\}$ is clearly not sum-free. Nonetheless, as recalled in Section 2.3, $\mathfrak{c}_{2i+1}(X,Y)=I_{\{0\}\cup\delta_{2i}}$ is an atom for every $i\geq 3$. In this section, we will provide a further class of atoms not accounted for by Theorem 3.3.
@@
Before proving this result, we need some technical lemmas based on the properties of the positive integers $a_{1},\dots,a_{n+1}$
@@
Then, $I\cup J=\llbracket 1,n-1\rrbracket$ and $H=I\cap J$.
@@
Moreover $n-1\in H$ if and only if $n-1\in I\cap J$. In fact if $n-1\in H$ and $n-1\notin I\cap J$, then
@@
On the other hand, let $I\subseteq\llbracket 1,n-1\rrbracket$, $J\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\cup\{Y\cup\{n+1\}\colon Y\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\}$ and assume $a_{I}+a_{J}\in B_{n}$. Then $a_{I}+a_{J}=a_{H}$, where either $H\subseteq\llbracket 1,n-1\rrbracket$, or $H=\llbracket 1,n\rrbracket$ or $H=X\cup\{n+1\}$, with $X\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$. Let us first observe that $H\neq\llbracket 1,n\rrbracket$. Indeed, if we assume that $H=\llbracket 1,n\rrbracket$, if $n+1\in J$, then $a_{I}+a_{J}\geq a_{n+1}>a_{H}$, while if $n+1\notin J$ (so $I,J\subseteq\llbracket 1,n-1\rrbracket$), $a_{I}+a_{J}\leq 2a_{\llbracket 1,n-1\rrbracket}<a_{n}\leq a_{H}$.
@@
Finally, we prove that $a_{\llbracket 1,n\rrbracket}+a_{I}\in B_{n}$ only if $I=\emptyset$ (the reverse implication is trivial). Assume $a_{\llbracket 1,n\rrbracket}+a_{I}=a_{H}$, where either $H\subseteq\llbracket 1,n-1\rrbracket$, $H=\llbracket 1,n\rrbracket$ or $H=X\cup\{n+1\}$, with $X\subseteq\llbracket 1,n-1\rrbracket$. First observe that $H$ cannot contain $n+1$, otherwise $a_{H}\geq a_{n+1}=a_{\llbracket 1,n\rrbracket}+a_{n}>a_{\llbracket 1,n\rrbracket}+2a_{\llbracket 1,n-1\rrbracket}\geq a_{\llbracket 1,n\rrbracket}+a_{I}$, so we are left to check only the first two of the above conditions. If $H\subseteq\llbracket 1,n-1\rrbracket$, Lemma 3.8 implies $\llbracket 1,n\rrbracket\cap I=\emptyset$ and this can happen only if $I=\emptyset$. If $H=\llbracket 1,n\rrbracket$, then $a_{I}=0$ and $I=\emptyset$.
(v) Firstly, if $I\cap J=\emptyset$, then $a_{I}+a_{J}=a_{I\cup J}\in C_{n}$, while if $I\cup J=\llbracket 1,n\rrbracket$ and $n\in I\cap J$, then $a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}=a_{\llbracket 1,n\rrbracket}+a_{n}+a_{((I\cap J)\setminus\{n\})}=a_{((I\cap J)\setminus\{n\})\cup\{n+1\}}\in C_{n}$.
@@
If $h=h^{\prime}$, we distinguish two sub-cases. If $h=h^{\prime}\leq n$, then $I,J,H\subseteq\llbracket 1,n\rrbracket$ and we get from Lemma 3.8 (with $k=n$) and (6) that $I\cap J=\emptyset$. If $h=h^{\prime}=n+1$, then $n+1\notin I\cap J$ (otherwise $a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}\geq 2a_{n+1}>a_{H}$). Thus, (6) becomes $a_{H\setminus\{n+1\}}=a_{(I\cup J)\setminus\{n+1\}}+a_{I\cap J}$ and we get from Lemma 3.8 with $k=n$, that $(I\cup J)\setminus\{n+1\}$ and $I\cap J$ have no elements in common, i.e., $I\cap J=\emptyset$.
It remains to consider the case $h>h^{\prime}$. This can occur only if $h=n+1$ and $h^{\prime}=n$. In fact, if $h=n+1$ and $h^{\prime}\leq n-1$, then $a_{H}\geq a_{n+1}>a_{n}>2a_{\llbracket 1,n-1\rrbracket}\geq a_{I}+a_{J}$, while if $h\leq n$, then $a_{H}\geq a_{h}>2a_{\llbracket 1,h-1\rrbracket}\geq a_{I}+a_{J}$. Moreover, it must be $n\in I\cap J$, otherwise we get the following contradiction: $a_{H}\geq a_{n+1}>2a_{n}>a_{n}+2a_{\llbracket 1,n-1\rrbracket}\geq a_{I}+a_{J}$. In addition, we have $n\notin H$, otherwise $a_{H}\geq a_{n}+a_{n+1}=2a_{n}+a_{\llbracket 1,n\rrbracket}>2a_{n}+3a_{\llbracket 1,n-1\rrbracket}\geq 2a_{\llbracket 1,n\rrbracket}\geq a_{I}+a_{J}$, which is again an absurd. Finally, summing up all the observations, we have $\max H=n+1$, $\max I\cup J=n$, $H\setminus\{n+1\}\subseteq\llbracket 1,n-1\rrbracket$, $I\cup J\subseteq\llbracket 1,n\rrbracket$ and $n\in I\cap J$. Then, (6) becomes $a_{H}=a_{H\setminus\{n+1\}}+a_{n+1}=a_{H\setminus\{n+1\}}+a_{\llbracket 1,n-1\rrbracket}+2a_{n}=a_{I\setminus\{n\}}+a_{J\setminus\{n\}}+2a_{n}=a_{I}+a_{J}$, and we get
@@
It then follows from Lemma 3.9 that $(I\setminus\{n\})\cup(J\setminus\{n\})=\llbracket 1,n-1\rrbracket$, thus $I\cup J=\llbracket 1,n\rrbracket$ and we already know that $n\in I\cap J$. All possible cases have been examined. ∎
@@
Set $\mathcal{P}:=\mathcal{P}(\llbracket 1,n-1\rrbracket)$, $\mathcal{P}^{+}:=\mathcal{P}\setminus\{\emptyset\}$, $\mathcal{Q}:=\mathcal{P}\cup\{Y\cup\{n+1\}\colon Y\in\mathcal{P}\}$ and $\mathcal{Q}^{+}:=\mathcal{Q}\setminus\{\emptyset\}$. For every fixed $i\in\llbracket 2,r\rrbracket$, $f_{i}g_{1}$ belongs to $(I_{B_{n}})_{K}[m]$, whence also $\operatorname{\mathsf{in}}(f_{i}g_{1})=X^{m-\alpha_{i}}Y^{\alpha_{i}}=\in(I_{B_{n}})_{K}[m]$. Since $1\leq\alpha_{i}\leq d\leq a_{\llbracket 1,n\rrbracket}$, by (7) it follows that $\alpha_{i}\in B_{n}\cap\llbracket 1,a_{\llbracket 1,n\rrbracket}\rrbracket=A_{n}\cup\{a_{\llbracket 1,n\rrbracket}\}\setminus{\{0\}}$, so we can write $\alpha_{i}=a_{I}$ for some $I\in\mathcal{P}^{+}\cup\{\llbracket 1,n\rrbracket\}$. Likewise, by looking at $\operatorname{\mathsf{in}}(f_{1}g_{j})$ with $j\in\llbracket 2,s\rrbracket$, we conclude that $\beta_{j}\in B_{n}\setminus{\{0,m\}}$, and so we can write $\beta_{j}=a_{J}$ for some $J\in\mathcal{Q}^{+}\cup\{\llbracket 1,n\rrbracket\}\setminus{\{\llbracket 1,n-1\rrbracket\cup\{n+1\}\}}$. Thus, $\operatorname{\mathsf{in}}(f_{i}g_{j})=X^{m-(a_{I}+a_{J})}Y^{a_{I}+a_{J}}$ and by (7) we conclude $a_{I}+a_{J}\in B_{n}\setminus{\{0\}}$.
Assume that for some $i\in\llbracket 2,r\rrbracket$, $\operatorname{\mathsf{in}}(f_{i})=Y^{a_{\llbracket 1,n\rrbracket}}$, i.e. $I=\llbracket 1,n\rrbracket$. This can only happen if $d=e=m/2=a_{\llbracket 1,n\rrbracket}$. For some $j\in\llbracket 2,s\rrbracket$, consider $\operatorname{\mathsf{in}}(g_{j})=X^{e-a_{J}}Y^{a_{J}}$, with $a_{J}\geq e$. If $a_{J}<e=a_{\llbracket 1,n\rrbracket}$, then $J\in\mathcal{P}^{+}$ and, by Lemma 3.10(iv), $a_{J}+a_{\llbracket 1,n\rrbracket}\notin B_{n}$, implying $f_{i}g_{j}\notin(I_{B_{n}})_{K}[m]$. Hence, the only possibility is $a_{J}=e$, whence we must have $\dim\mathfrak{b}_{K}[e]=2$. An analogous argument shows that also $\dim\mathfrak{a}_{K}[d]=2$, and so $\dim(\mathfrak{a}_{K}[d]\cdot\mathfrak{b}_{K}[e])\leq 4$, contradicting (7). In fact $\dim(I_{B_{n}})_{K}[m]=|B_{n}|=2^{n}+1\geq 5$.
@@
(2) $\{2,n+1\}\subseteq\mathsf{L}_{\mathcal{I}(R)}(I_{C_{n}})$ follows by Remark 3.6, the previous point (1) and the fact that all the ideals of the form $\mathfrak{b}_{i}$ are atoms of ${\mathcal{I}}(R)$ (by [21, Prop. 5.10.1]).
@@
shows that $a_{ij}\in C_{n}$ and hence $a_{ij}=a_{I_{j}^{(i)}}$ with $\emptyset\neq I_{j}^{(i)}\subseteq\llbracket 1,n+1\rrbracket$ for every $i\in\llbracket 1,n+2\rrbracket$ and $j\in\llbracket 2,r_{i}\rrbracket$. Likewise, $\sum_{\ell=1}^{k}a_{I_{j_{\ell}}^{(i_{\ell})}}\in C_{n}$ for every $k\leq n+2$ and $i_{s}\neq i_{t}$ for every $s\neq t$. We claim that $n\in I_{2}^{(i)}$ for all $i\in\llbracket 1,n+2\rrbracket$ and that $I_{2}^{(i)}\subseteq\llbracket 1,n\rrbracket$ for all $i\in\llbracket 1,n+2\rrbracket$. If $n\not\in I_{2}^{(i_{0})}$ for some $i_{0}$, since $a_{I_{2}^{(i_{0})}}+a_{I_{2}^{(i)}}\in C_{n}$ for every $i\neq i_{0}$, by Lemma 3.10(v) we get
@@
In [16, Prop. 4.10], the authors show that $B_{n}$ is an atom in $\mathcal{P}_{\text{fin},0}(\mathbb{N})$ and that $\mathsf{L}_{\mathcal{P}_{\text{fin},0}(\mathbb{N})}(I_{C_{n}})=\{2,n+1\}$. In light of theorem 3.7, it makes sense to ask whether, in analogy with the results in [16], also $\mathsf{L}_{\mathcal{I}(R)}(I_{C_{n}})=\{2,n+1\}$; we note that this would immediately imply $\Delta({\mathcal{I}}(R))=\mathbb{N}^{+}$. The question remains open but in the next section we will prove that in the case $D=K$ the answer is negative, if we restrict the problem from $\mathcal{I}(R)$ to its submonoid ${\mathcal{M}{\rm on}}(R)$ of nonzero monomial ideals in $R$.

### v2

@@ 3.1. Atoms constructed from sum-free sets of positive integers
We are now in the position to extend the results on the atomicity of $\mathfrak{b}_{i}(X,Y)$ and $\mathfrak{c}_{2i+1}(X,Y)$ to a larger family of ideals. Before proceeding, let us recall a classical definition from additive combinatorics: a nonempty subset $A\subseteq\mathbb{N}$ is sum-free if $A\cap(A+A)=\emptyset$. Easy examples of sum-free sets include any set of odd positive integers and any subset of $\{\lceil{(n+1)/2}\rceil,\dots,n\}$, where $n\in\mathbb{N}$ and $\lceil\ \rceil$ denotes the usual ceiling function. In fact, Green proved in [28] that these examples already give the correct order of magnitude for the number of sum-free sets: more precisely, there are $O(2^{\nicefrac{{n}}{{2}}})$ sum-free subsets of $\{1,\dots,n\}$ for $n\to\infty$, as predicted by the Cameron-Erdös Conjecture.
@@
For every $i\in\mathbb{N}^{+}$, the sets $\beta_{i}=\{i\}$ and $\delta_{2i+1}=\{2\llbracket 0,i\rrbracket+1\}$ are clearly sum-free. We then get from Lemma 3.2 an alternative proof of the fact (mentioned in the introduction and proved in [16]) that $\{0\}\cup\beta_{i}$ and $\{0\}\cup\delta_{2i+1}$ are atoms of $\mathcal{P}_{\rm fin}(\mathbb{N})$. The next theorem (the first of the main results of this section) provides a general method for constructing atoms in ${\mathcal{I}}(R)$ from sum-free subsets of $\mathbb{N}$ and, as a special case, yields the atomicity of $\mathfrak{b}_{i}(X,Y)=I_{\{0\}\cup\beta_{i}}$ and $\mathfrak{c}_{2i+1}(X,Y)=I_{\{0\}\cup\delta_{2i+1}}$.
@@ 3.2. A further family of atoms
For every $i\in\mathbb{N}_{\geq 2}$, the set $\delta_{2i}=\{1,2\llbracket 1,i\rrbracket\}$ is clearly not sum-free. Nonetheless, as recalled in Section 2.3, $\mathfrak{c}_{2i+1}(X,Y)=I_{\{0\}\cup\delta_{2i}}$ is an atom for every $i\geq 3$. In this section, we will provide a further class of atoms not covered by Theorem 3.3.
@@
Before proving this result, we need some technical lemmas based on the properties of the positive integers $a_{1},\dots,a_{n+1}$.
@@
Then $I\cup J=\llbracket 1,n-1\rrbracket$ and $H=I\cap J$.
@@
Moreover $n-1\in H$ if and only if $n-1\in I\cap J$. In fact, if $n-1\in H$ and $n-1\notin I\cap J$, then
@@
On the other hand, let $I\subseteq\llbracket 1,n-1\rrbracket$, $J\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\cup\{Y\cup\{n+1\}\colon Y\in\mathcal{P}(\llbracket 1,n-1\rrbracket)\}$ and assume $a_{I}+a_{J}\in B_{n}$. Then $a_{I}+a_{J}=a_{H}$, where either $H\subseteq\llbracket 1,n-1\rrbracket$, $H=\llbracket 1,n\rrbracket$ or $H=X\cup\{n+1\}$ with $X\in\mathcal{P}(\llbracket 1,n-1\rrbracket)$. Let us first observe that $H\neq\llbracket 1,n\rrbracket$. Indeed, if we assume that $H=\llbracket 1,n\rrbracket$, if $n+1\in J$, then $a_{I}+a_{J}\geq a_{n+1}>a_{H}$, while if $n+1\notin J$ (so $I,J\subseteq\llbracket 1,n-1\rrbracket$), $a_{I}+a_{J}\leq 2a_{\llbracket 1,n-1\rrbracket}<a_{n}\leq a_{H}$.
@@
Finally, we prove that $a_{\llbracket 1,n\rrbracket}+a_{I}\in B_{n}$ only if $I=\emptyset$ (the reverse implication is trivial). Assume $a_{\llbracket 1,n\rrbracket}+a_{I}=a_{H}$, where either $H\subseteq\llbracket 1,n-1\rrbracket$, $H=\llbracket 1,n\rrbracket$ or $H=X\cup\{n+1\}$ with $X\subseteq\llbracket 1,n-1\rrbracket$. First, observe that $H$ cannot contain $n+1$, otherwise $a_{H}\geq a_{n+1}=a_{\llbracket 1,n\rrbracket}+a_{n}>a_{\llbracket 1,n\rrbracket}+2a_{\llbracket 1,n-1\rrbracket}\geq a_{\llbracket 1,n\rrbracket}+a_{I}$, so we are left to check only the first two of the above conditions. If $H\subseteq\llbracket 1,n-1\rrbracket$, Lemma 3.8 implies $\llbracket 1,n\rrbracket\cap I=\emptyset$ and this can happen only if $I=\emptyset$. If $H=\llbracket 1,n\rrbracket$, then $a_{I}=0$ and $I=\emptyset$.
(v) First, if $I\cap J=\emptyset$, then $a_{I}+a_{J}=a_{I\cup J}\in C_{n}$, while if $I\cup J=\llbracket 1,n\rrbracket$ and $n\in I\cap J$, then $a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}=a_{\llbracket 1,n\rrbracket}+a_{n}+a_{((I\cap J)\setminus\{n\})}=a_{((I\cap J)\setminus\{n\})\cup\{n+1\}}\in C_{n}$.
@@
If $h=h^{\prime}$, we distinguish two sub-cases. If $h=h^{\prime}\leq n$, then $I,J,H\subseteq\llbracket 1,n\rrbracket$ and we get from Lemma 3.8 (with $k=n$) and (6) that $I\cap J=\emptyset$. If $h=h^{\prime}=n+1$, then $n+1\notin I\cap J$ (otherwise, $a_{I}+a_{J}=a_{I\cup J}+a_{I\cap J}\geq 2a_{n+1}>a_{H}$). Thus, (6) becomes $a_{H\setminus\{n+1\}}=a_{(I\cup J)\setminus\{n+1\}}+a_{I\cap J}$ and we get from Lemma 3.8 with $k=n$, that $(I\cup J)\setminus\{n+1\}$ and $I\cap J$ have no elements in common, i.e., $I\cap J=\emptyset$.
It remains to consider the case $h>h^{\prime}$. This can occur only if $h=n+1$ and $h^{\prime}=n$. In fact, if $h=n+1$ and $h^{\prime}\leq n-1$, then $a_{H}\geq a_{n+1}>a_{n}>2a_{\llbracket 1,n-1\rrbracket}\geq a_{I}+a_{J}$, while if $h\leq n$, then $a_{H}\geq a_{h}>2a_{\llbracket 1,h-1\rrbracket}\geq a_{I}+a_{J}$. Moreover, it must be $n\in I\cap J$, otherwise we get the following contradiction: $a_{H}\geq a_{n+1}>2a_{n}>a_{n}+2a_{\llbracket 1,n-1\rrbracket}\geq a_{I}+a_{J}$. In addition, we have $n\notin H$, otherwise $a_{H}\geq a_{n}+a_{n+1}=2a_{n}+a_{\llbracket 1,n\rrbracket}>2a_{n}+3a_{\llbracket 1,n-1\rrbracket}\geq 2a_{\llbracket 1,n\rrbracket}\geq a_{I}+a_{J}$, which is again absurd. Finally, summing up all the observations, we have $\max H=n+1$, $\max I\cup J=n$, $H\setminus\{n+1\}\subseteq\llbracket 1,n-1\rrbracket$, $I\cup J\subseteq\llbracket 1,n\rrbracket$ and $n\in I\cap J$. Then (6) becomes $a_{H}=a_{H\setminus\{n+1\}}+a_{n+1}=a_{H\setminus\{n+1\}}+a_{\llbracket 1,n-1\rrbracket}+2a_{n}=a_{I\setminus\{n\}}+a_{J\setminus\{n\}}+2a_{n}=a_{I}+a_{J}$, and we get
@@
It follows from Lemma 3.9 that $(I\setminus\{n\})\cup(J\setminus\{n\})=\llbracket 1,n-1\rrbracket$, thus $I\cup J=\llbracket 1,n\rrbracket$ and we already know that $n\in I\cap J$. All possible cases have been examined. ∎
@@
Set $\mathcal{P}:=\mathcal{P}(\llbracket 1,n-1\rrbracket)$, $\mathcal{P}^{+}:=\mathcal{P}\setminus\{\emptyset\}$, $\mathcal{Q}:=\mathcal{P}\cup\{Y\cup\{n+1\}\colon Y\in\mathcal{P}\}$ and $\mathcal{Q}^{+}:=\mathcal{Q}\setminus\{\emptyset\}$. For every fixed $i\in\llbracket 2,r\rrbracket$, $f_{i}g_{1}$ belongs to $(I_{B_{n}})_{K}[m]$, whence also $\operatorname{\mathsf{in}}(f_{i}g_{1})=X^{m-\alpha_{i}}Y^{\alpha_{i}}=\in(I_{B_{n}})_{K}[m]$. Since $1\leq\alpha_{i}\leq d\leq a_{\llbracket 1,n\rrbracket}$, by (7) it follows that $\alpha_{i}\in B_{n}\cap\llbracket 1,a_{\llbracket 1,n\rrbracket}\rrbracket=A_{n}\cup\{a_{\llbracket 1,n\rrbracket}\}\setminus{\{0\}}$, so we can write $\alpha_{i}=a_{I}$ for some $I\in\mathcal{P}^{+}\cup\{\llbracket 1,n\rrbracket\}$. Similarly, by looking at $\operatorname{\mathsf{in}}(f_{1}g_{j})$ with $j\in\llbracket 2,s\rrbracket$, we conclude that $\beta_{j}\in B_{n}\setminus{\{0,m\}}$, so we can write $\beta_{j}=a_{J}$ for some $J\in\mathcal{Q}^{+}\cup\{\llbracket 1,n\rrbracket\}\setminus{\{\llbracket 1,n-1\rrbracket\cup\{n+1\}\}}$. Thus, $\operatorname{\mathsf{in}}(f_{i}g_{j})=X^{m-(a_{I}+a_{J})}Y^{a_{I}+a_{J}}$ and by (7) we conclude $a_{I}+a_{J}\in B_{n}\setminus{\{0\}}$.
Assume that for some $i\in\llbracket 2,r\rrbracket$, $\operatorname{\mathsf{in}}(f_{i})=Y^{a_{\llbracket 1,n\rrbracket}}$, i.e. $I=\llbracket 1,n\rrbracket$. This can only happen if $d=e=m/2=a_{\llbracket 1,n\rrbracket}$. For some $j\in\llbracket 2,s\rrbracket$, consider $\operatorname{\mathsf{in}}(g_{j})=X^{e-a_{J}}Y^{a_{J}}$, with $a_{J}\geq e$. If $a_{J}<e=a_{\llbracket 1,n\rrbracket}$, then $J\in\mathcal{P}^{+}$ and, by Lemma 3.10(iv), $a_{J}+a_{\llbracket 1,n\rrbracket}\notin B_{n}$, implying $f_{i}g_{j}\notin(I_{B_{n}})_{K}[m]$. Hence, the only possibility is $a_{J}=e$, so we must have $\dim\mathfrak{b}_{K}[e]=2$. An analogous argument shows that also $\dim\mathfrak{a}_{K}[d]=2$, and so $\dim(\mathfrak{a}_{K}[d]\cdot\mathfrak{b}_{K}[e])\leq 4$, contradicting (7). In fact $\dim(I_{B_{n}})_{K}[m]=|B_{n}|=2^{n}+1\geq 5$.
@@
(2) $\{2,n+1\}\subseteq\mathsf{L}_{{\mathcal{I}}(R)}(I_{C_{n}})$ follows by Remark 3.6, the previous point (1) and the fact that all ideals of the form $\mathfrak{b}_{i}$ are atoms of ${\mathcal{I}}(R)$ (by [21, Prop. 5.10.1]).
@@
shows that $a_{ij}\in C_{n}$ and hence $a_{ij}=a_{I_{j}^{(i)}}$ with $\emptyset\neq I_{j}^{(i)}\subseteq\llbracket 1,n+1\rrbracket$ for every $i\in\llbracket 1,n+2\rrbracket$ and $j\in\llbracket 2,r_{i}\rrbracket$. Similarly, $\sum_{\ell=1}^{k}a_{I_{j_{\ell}}^{(i_{\ell})}}\in C_{n}$ for every $k\leq n+2$ and $i_{s}\neq i_{t}$ for every $s\neq t$. We claim that $n\in I_{2}^{(i)}$ for all $i\in\llbracket 1,n+2\rrbracket$ and that $I_{2}^{(i)}\subseteq\llbracket 1,n\rrbracket$ for all $i\in\llbracket 1,n+2\rrbracket$. If $n\not\in I_{2}^{(i_{0})}$ for some $i_{0}$, since $a_{I_{2}^{(i_{0})}}+a_{I_{2}^{(i)}}\in C_{n}$ for every $i\neq i_{0}$, by Lemma 3.10(v) we get
@@
In [16, Prop. 4.10], the authors show that $B_{n}$ is an atom in $\mathcal{P}_{\text{fin},0}(\mathbb{N})$ and that $\mathsf{L}_{\mathcal{P}_{\text{fin},0}(\mathbb{N})}(C_{n})=\{2,n+1\}$. In light of theorem 3.7, it makes sense to ask whether, in analogy with the results in [16], also $\mathsf{L}_{{\mathcal{I}}(R)}(I_{C_{n}})=\{2,n+1\}$; we note that this would immediately imply $\Delta({\mathcal{I}}(R))=\mathbb{N}^{+}$. The question remains open, but in the next section we will prove that in the case $D=K$ the answer is negative, if we consider the analogous problem in the submonoid ${\mathcal{M}{\rm on}}(R)\subseteq{\mathcal{I}}(R)$ of nonzero monomial ideals in $R$.